# On the multiplier spectrum of polynomials

### v1

@@ Abstract.
We prove several results on multiplier spectrum for polynomials. We provide a detailed proof of the theorem stating that the multiplier spectrum morphism is generically injective on the moduli space of polynomials. We obtain a description of the non-injective locus of the multiplier spectrum morphism for polynomials of all degrees $d\geq 2$. Roughly speaking, non-injectivity implies intertwining, as it signifies the equivalence of polynomials or Ritt moves, except at isolated points and up to iteration. We investigate the relation between Ritt moves and the multiplier spectrum over arithmetic progressions.
@@ 1. Introduction
The aim of this paper is to study the multipliers of periodic points for a rational map $f$ on $\mathbb{P}^{1}(\mathbb{C})$ of degree at least $2$, concentrating on the case when $f$ is a polynomial.
@@ 1.1. The multiplier spectrum morphism on rational maps
Let $f:\mathbb{P}^{1}\to\mathbb{P}^{1}$ be a rational map over $\mathbb{C}$ of degree $d\geq 2$. For a $f$-periodic point $z_{0}\in\mathbb{P}^{1}(\mathbb{C})$ with exact period $n$, we set $n_{f}(z_{0}):=n$ to denote this exact period. The multiplier $\rho_{f}(z_{0})$ of $f$ at $z_{0}$ is defined as the differential $df^{\circ n}(z_{0})\in\mathbb{C}$. (Throughout this thesis, the symbol $\circ$ is always used to denote composition of rational maps, while its absence indicates multiplication between rational maps.) We write $n(z_{0})=n_{f}(z_{0})$ and $\rho(z_{0})=\rho_{f}(z_{0})$ for simplicity, when the map $f$ is clear. The length of $f$ at $z_{0}$ is defined as the modulus $\lvert\rho_{f}(z_{0})\rvert$. The characteristic exponent of $f$ at $z_{0}$ is defined as $\chi_{f}(z_{0}):=n^{-1}\log\lvert\rho_{f}(z_{0})\rvert$ (when $\rho_{f}(z_{0})\neq 0$). Clearly, the multiplier, length, and characteristic exponent are all invariant under conjugacy by Möbius transformations.
Let ${\rm Per}(f)(\mathbb{C})$ denote the set of all periodic points in $\mathbb{P}^{1}(\mathbb{C})$ of $f$, and define
${\rm Per}^{*}(f)(\mathbb{C}):=\{z_{0}\in{\rm Per}(f)(\mathbb{C}):\rho_{f}(z_{0})\neq 0\}.$
We write ${\rm Per}(f)$ and ${\rm Per}^{*}(f)$ for simplicity when the base field $\mathbb{C}$ is clear. We let ${\rm PrePer}(f)(\mathbb{C})$ (or ${\rm PrePer}(f)$) denote the set of all $f$-preperiodic points in $\mathbb{P}^{1}(\mathbb{C})$.
We identify ${\rm PGL}_{2}(\mathbb{C})$ with the group of rational maps $g\in\mathbb{C}(z)$ of degree $1$ (i.e., the automorphism group of $\mathbb{P}^{1}$ over $\mathbb{C}$).
For an endomorphism $g$ on an algebraic variety $X$ and an endomorphism $h$ on an algebraic variety $Y$, we say $h$ is semi-conjugate to $g$ if there is a dominant morphism $\pi:X\to Y$ such that $\pi\circ g=h\circ\pi$, and we write $g\geq h$ (or $g\geq_{\pi}h$ when $\pi$ is specified).
In complex dynamics, the exceptional maps defined below are often regarded as special examples among all rational maps of degree $\geq 2$. These are rational maps related to algebraic groups that exhibit special dynamical properties.
@@
  • The map $f$ is called Lattès if there exists an endomorphism $\phi$ on an elliptic curve $E$ such that $\phi\geq f$. Furthermore, the map $f$ is called flexible Lattès if there exist an elliptic curve $E$ and $n\in\mathbb{Z}\setminus\{0,\pm 1\}$ such that $n]\geq_{\pi}f$, where $[n]$ denotes the multiplication-by-$n$ map on $E$ and $\pi:E\to\mathbb{P}^{1}$ is the quotient map modulo $\{\pm 1\}$. (Note that in this case we must have $d=n^{2}$, a perfect square.) A non-flexible Lattès map is called rigid Lattès.
  • We say that $f$ is of monomial type if it is semi-conjugate to the power map $z\mapsto z^{n}$ on $\mathbb{P}^{1}$ for some integer $n\neq 0,\pm 1$. It is known that $f$ is of monomial type if and only if it is conjugate to $z^{\pm d}$ or $\pm T_{d}(z)$, where $T_{d}(z)$ is the Chebyshev polynomial of degree $d$; see [Mil06].
@@
These definitions depend only on the conjugacy class of $f$. It is well-known that $f$ is exceptional if and only if some (hence every) iterate $f^{\circ k}$ is exceptional ($k\in\mathbb{Z}_{>0}$).
Fix an integer $d\geq 2$. Let ${\rm Rat}_{d}$ be the space of degree $d$ endomorphisms on $\mathbb{P}^{1}$, which is a smooth affine variety of dimension $2d+1$ [Sil12]. Let ${\rm FL}_{d}\subseteq{\rm Rat}_{d}$ be the locus of flexible Lattès maps, which is Zariski closed in ${\rm Rat}_{d}$. The group ${\rm PGL}_{2}={\rm Aut}(\mathbb{P}^{1})$ acts on ${\rm Rat}_{d}$ by conjugation. The geometric quotient
$\mathcal{M}_{d}:={\rm Rat}_{d}/{\rm PGL}_{2}={\rm Spec}(\mathcal{O}({\rm Rat}_{d})^{{\rm PGL}_{2}})$
is the (coarse) moduli space of endomorphisms of degree $d$ [Sil12], which is an affine variety of dimension 2d22d-2 [Sil07, Theorem 4.36(c)]. Let $\Psi:{\rm Rat}_{d}\to\mathcal{M}_{d}$ be the quotient morphism. For $f\in{\rm Rat}_{d}(\mathbb{C})$, we also denote its conjugacy class by $\Psi(f)=[f]$. Set ${\rm FL}_{d}]:=\Psi({\rm FL}_{d})$. Note that ${\rm FL}_{d}$ is non-empty if and only if $d$ is a perfect square, and in this case ${\rm FL}_{d}]$ is equidimensional of dimension $1$. The above construction works over any algebraically closed field of characteristic 0 and commutes with base changes.
Fix $n\in\mathbb{Z}_{>0}$. Let ${\rm Per}_{n}(f)={\rm Per}_{n}(f)(\mathbb{C})$ be the multi-set [^1] A multi-set is a set except allowing multiple instances for each of its elements. The number of the instances of an element is called its multiplicity. For example, $\{a,a,b,c,c,c\}$ is a multi-set of cardinality $6$, with multiplicities $2,1,3$ for $a,b,c$, respectively. of fixed points of $f^{\circ n}$. This multi-set has cardinality $d^{n}+1=:N_{d,n}$. Using elementary symmetric polynomials, the multipliers of the fixed points of $f^{\circ n}$ define a point
@@
where $\sigma_{j,n}(f)$ is the $j$-th elementary symmetric polynomial in the $N_{d,n}$-tuple $\{\rho_{f^{\circ n}}(z)\}_{z\in{\rm Per}_{n}(f)}$ for $1\leq j\leq N_{d,n}$. We define the multiplier spectrum of $f$ to be the sequence $S(f)=(S_{n}(f))_{n=1}^{\infty}$. It is clear that $S_{n}$ takes the same value within a conjugacy class of rational maps and thus defines a morphism $S_{n}:\mathcal{M}_{d}(\mathbb{C})\to\mathbb{A}^{N_{d,n}}(\mathbb{C})$, called the multiplier spectrum morphism of level $n$. We define the multiplier spectrum morphism up to level $n$ as
$\tau_{d,n}:\mathcal{M}_{d}(\mathbb{C})\to\mathbb{A}^{N_{d,1}}(\mathbb{C})\times\cdots\times\mathbb{A}^{N_{d,n}}(\mathbb{C}),[f]\mapsto(S_{1}([f]),\dots,S_{n}([f])).$
For $f\in{\rm Rat}_{d}(\mathbb{C})$, we also write $\tau_{d,n}(f):=\tau_{d,n}([f])$.
For $n\in\mathbb{Z}_{>0}$, set $R_{d,n}:=\{([f],[g])\in\mathcal{M}_{d}(\mathbb{C})^{2}\mid\tau_{d,n}([f])=\tau_{d,n}([g])\}$. Then $(R_{d,n})_{n=1}^{\infty}$ forms a decreasing sequence of Zariski closed subsets of $\mathcal{M}_{d}(\mathbb{C})^{2}$, which stabilizes by noetherianity. Let $m_{d}\geq 1$ be minimal such that $\cap_{n=1}^{\infty}R_{d,n}=R_{d,m_{d}}$. Consequently, for all $g\in{\rm Rat}_{d}(\mathbb{C}), ff$ and $g$ have the same multiplier spectrum if and only if $\tau_{d,m_{d}}(f)=\tau_{d,m_{d}}(g)$, i.e., they have the same multiplier spectrum up to level $m_{d}$. We define $\tau_{d}:=\tau_{d,m_{d}}$, called the multiplier spectrum morphism on $\mathcal{M}_{d}$ (or ${\rm Rat}_{d}$). It is well-known that rational maps within an irreducible component of ${\rm FL}_{d}(\mathbb{C})$ share the same multiplier spectrum.
@@
$\tau_{d}:\mathcal{M}_{d}(\mathbb{C})\setminus\Psi({\rm FL}_{d}(\mathbb{C}))\to\mathbb{A}^{N_{d,1}}(\mathbb{C})\times\cdots\times\mathbb{A}^{N_{d,m_{d}}}(\mathbb{C})$
@@
Similarly, by replacing the multiplier with its modulus and $\mathbb{C}$ with $\mathbb{R}$ in the above definitions, we obtain for each $n\in\mathbb{Z}_{>0}$ the length spectrum map $L_{n}:\mathcal{M}_{d}(\mathbb{C})\to\mathbb{R}^{N_{d,n}}$ of level $n$. The length spectrum of $f$ is defined as the sequence $L(f)=(L_{n}(f))_{n=1}^{\infty}$. We define the length spectrum map up to level $n$ as
$\eta_{d,n}:\mathcal{M}_{d}(\mathbb{C})\to\mathbb{R}^{N_{d,1}}\times\cdots\times\mathbb{R}^{N_{d,n}},[f]\mapsto(L_{1}([f]),\dots,L_{n}([f])).$
(For the time being, we treat these maps merely as maps between sets.)
@@ 1.2. Statement of the main results
Fix an integer $d\geq 2$. By restriction, we obtain the multiplier spectrum morphism for polynomials. Let ${\rm Poly}^{d}\subset{\rm Rat}_{d}$ be the closed subvariety of polynomials of degree $d$, and let ${\rm MPoly}^{d}:={\rm Poly}^{d}/{\rm Aff}$ be the moduli space of polynomials of degree $d$, see [FG22]. Here $\rm Aff$} is the group of automorphisms of $\mathbb{A}^{1}$ consisting of all linear polynomials, which acts on ${\rm Poly}^{d}$ by conjugation. It is clear that for all $g\in{\rm Poly}^{d}(\mathbb{C}), ff$ and $g$ are conjugate by a Möbius transformation $\tau\in{\rm PGL}_{2}(\mathbb{C})$ if and only if they are conjugate by a linear polynomial $\sigma\in{\rm Aff}(\mathbb{C})$. Let $\Psi:{\rm Poly}^{d}\to{\rm MPoly}^{d}$ still denote the quotient morphism. For every $f\in{\rm Poly}^{d}(\mathbb{C})$, we also let $\Psi(f)=[f]\in{\rm MPoly}^{d}(\mathbb{C})$ denote its conjugacy class. We can view ${\rm MPoly}^{d}\subset\mathcal{M}_{d}$, which is compatible with $\Psi$ and ${\rm Poly}^{d}\subset{\rm Rat}_{d}$. For every integer $n\geq 1$, denote the restriction of $\tau_{d,n}$ (resp. $\tau_{d}$) to the moduli space of polynomials ${\rm MPoly}^{d}(\mathbb{C})$ by $\tilde{\tau}_{d,n}$ (resp. $\tilde{\tau}_{d}$). Similarly, we obtain $\tilde{\eta}_{d,n}$ for the length spectrum.
@@
Ji and Xie also prove the following polynomial version of Theorem 1.3, which is also proved by Huguin [Hug24] independently using completely different methods.
@@
$\tilde{\tau}_{d}:{\rm MPoly}^{d}(\mathbb{C})\to\mathbb{A}^{N_{d,1}}(\mathbb{C})\times\cdots\times\mathbb{A}^{N_{d,m_{d}}}(\mathbb{C})$
@@
In [JX25], the proof of Theorem 1.4 is similar to the proof of Theorem 1.3 in the same paper. Specifically, the proof of Theorem 1.3 relies on two key ingredients: a variant of the Dynamical André-Oort (DAO) conjecture for curves [JX25, Theorem 3.4] and a result [JX25, Theorem 3.3] based on Pakovich’s work [Pak25a]. In contrast, the proof of Theorem 1.4 replaces the latter ingredient with [FG22, Theorems 3.51 and 3.52]. Huguin’s proof [Hug24] actually shows that $\tilde{\tau}_{d,2}$ is generically injective for every $d\geq 2$, and it is based on intricate local computations.
The result [JX25, Theorem 3.3] involves a specific class of rational maps, namely simple rational maps. Recall that a rational map $G\in\mathbb{C}(z)$ of degree $m\geq 2$ is called simple if $G$ has exactly 2m22m-2 distinct critical values in $\mathbb{P}^{1}(\mathbb{C})$. However, a polynomial of degree $d\geq 3$ is never simple, due to the totally invariant fixed point $\infty$. To overcome this problem, for a polynomial $F\in\mathbb{C}[z]$ of degree $m\geq 2$, we call $F$ pre-simple if it has exactly m1m-1 distinct critical values in $\mathbb{C}$. This class of pre-simple polynomials serves as an analogue of simple rational maps in the polynomial case. We find that Pakovich’s methods [Pak25a] are also applicable to pre-simple polynomials. In §3.1, we present the following polynomial version of [JX25, Theorem 3.3] (for the case $d\geq 4$), which is essentially due to Pakovich [Pak25a]:
@@
For every integer $d\geq 4$, there exists a non-empty Zariski open subset $U$ of ${\rm Poly}^{d}(\mathbb{C})$ such that, for every $f\in U$, the following hold:
@@
    for every pre-simple $g\in{\rm Poly}^{d}(\mathbb{C})$, if $f$ and $g$ are intertwined, then $[f]=[g]$ in ${\rm MPoly}^{d}(\mathbb{C})$.
Recall that two polynomials $F,G\in\mathbb{C}[z]$ of degree $\geq 2$ are called intertwined if there exists a (possibly reducible) algebraic curve $Z\subset(\mathbb{A}^{1})^{2}$ whose projections to both axes are onto, and $Z$ is invariant under the endomorphism $(F,G):(\mathbb{A}^{1})^{2}\to(\mathbb{A}^{1})^{2}$.
Recently, Pakovich [Pak25b] has studied general rational maps of degree $2\leq d\leq 3$ in detail. Consequently, one can prove Theorem 1.3 for this case using the results in [Pak25b] as part of the argument. The philosophy is that a general rational map of degree $d$ has “good” properties regarding the decompositions of its iterates. However, the simplicity condition is not sufficient when $2\leq d\leq 3$, unlike the situation in [Pak25a] where $d\geq 4$.
Following Pakovich [Pak25b], we study general polynomial maps of degree $2\leq d\leq 3$ in §3.2 and show the following polynomial version of [JX25, Theorem 3.3] for the case where $2\leq d\leq 3$:

### v2

@@ Abstract.
We prove several results on the multiplier spectrum of polynomials. We provide a detailed proof of the theorem stating that the multiplier spectrum morphism is generically injective on the moduli space of polynomials. We obtain a description of the non-injective locus of the multiplier spectrum morphism for polynomials of degree $d\geq 2$. Roughly speaking, we prove that, apart from isolated exceptions, polynomials with the same multiplier spectrum are intertwined. More precisely, we show that, up to iteration and isolated exceptions, the polynomials are either equivalent or related by Ritt moves. We also investigate the relationship between Ritt moves and multiplier spectra over arithmetic progressions.
@@ 1. Introduction
This paper aims to study the multipliers at periodic points of a rational map $f:\mathbb{P}^{1}(\mathbb{C})\to\mathbb{P}^{1}(\mathbb{C})$ of degree at least $2$, with a specific focus on the case where $f$ is a polynomial.
@@ 1.1. The multiplier spectrum morphism on rational maps
Let $f:\mathbb{P}^{1}\to\mathbb{P}^{1}$ be an endomorphism of degree $d\geq 2$ over $\mathbb{C}$, i.e., $f\in\mathbb{C}(z)$ is of degree $d\geq 2$. For an $f$-periodic point $z_{0}\in\mathbb{P}^{1}(\mathbb{C})$ with exact period $n$, set $n_{f}(z_{0}):=n$. The multiplier of $f$ at $z_{0}$ is defined as the differential
$\rho_{f}(z_{0}):=df^{\circ n}(z_{0})\in\mathbb{C}.$
(Throughout this paper, the symbol $\circ$ is always used to denote composition of rational maps, while its absence indicates multiplication in the field $\mathbb{C}(z)$ of rational maps. ) Write $n(z_{0})=n_{f}(z_{0})$ and $\rho(z_{0})=\rho_{f}(z_{0})$ for simplicity, when the map $f$ is clear from context. The length of $f$ at $z_{0}$ is the modulus $\left|\rho_{f}(z_{0})\right|$. The characteristic exponent of $f$ at $z_{0}$ is defined to be $\chi_{f}(z_{0}):=n^{-1}\log\left|\rho_{f}(z_{0})\right|$, when $\rho_{f}(z_{0})\neq 0$. Note that multipliers, lengths, and characteristic exponents are invariant under conjugacy by Möbius transformations.
Let $\mathrm{Per}(f)(\mathbb{C})$ denote the set of $f$-periodic points in $\mathbb{P}^{1}(\mathbb{C})$, and set
$\mathrm{Per}^{*}(f)(\mathbb{C}):=\{z_{0}\in\mathrm{Per}(f)(\mathbb{C})\colon\rho_{f}(z_{0})\neq 0\}.$
Write $\mathrm{Per}(f)$ and $\mathrm{Per}^{*}(f)$ for simplicity when the base field $\mathbb{C}$ is clear. Let $\mathrm{PrePer}(f)(\mathbb{C})$ (or $\mathrm{PrePer}(f)$) denote the set of $f$-preperiodic points in $\mathbb{P}^{1}(\mathbb{C})$.
We identify $\mathrm{PGL}_{2}(\mathbb{C})$ with the group of rational maps $g\in\mathbb{C}(z)$ of degree $1$ (i.e., Möbius transformations over $\mathbb{C}$), which forms the automorphism group of the projective line $\mathbb{P}^{1}$ over $\mathbb{C}$.
For endomorphisms $g:X\to X$ and $h:Y\to Y$ on algebraic varieties, we say that $h$ is semi-conjugate to $g$ if there is a dominant morphism $\pi:X\to Y$ such that $\pi\circ g=h\circ\pi$, written $g\geq h$ (or $g\geq_{\pi}h$ when $\pi$ is specified).
In complex dynamics, the exceptional maps defined below are often regarded as special examples within the class of rational maps of degree $\geq 2$. They are the rational maps on $\mathbb{P}^{1}(\mathbb{C})$ that are related to algebraic groups and exhibit special dynamical properties.
@@
  • The map $f$ is called a Lattès map if there exists an endomorphism $\phi$ on an elliptic curve $E$ such that $\phi\geq f$. The map $f$ is called flexible Lattès if there exist an elliptic curve $E$ and $n\in\mathbb{Z}\setminus\{0,\pm 1\}$ such that $n]\geq_{\pi}f$, where $[n]$ denotes the multiplication-by-$n$ map on $E$ and $\pi:E\to\mathbb{P}^{1}$ is the quotient map modulo $\{\pm 1\}$. (Note that in this case we must have $d=n^{2}$, a perfect square in $\mathbb{Z}$.) A non-flexible Lattès map is called rigid Lattès.
  • We say that $f$ is of monomial type if it is semi-conjugate to the power map $z\mapsto z^{n}$ on $\mathbb{P}^{1}$ for some $n\in\mathbb{Z}\setminus\{0,\pm 1\}$. It is well-known that $f$ is of monomial type if and only if it is conjugate to $z^{\pm d}$ or $\pm T_{d}(z)$, where $T_{d}(z)$ is the (normalized) Chebyshev polynomial of degree $d$ (or the Dickson polynomial of degree $d$ with parameter $1$), i.e., the unique monic polynomial $T_{d}\in\mathbb{C}[z]$ of degree $d$ such that $T_{d}(z+z^{-1})=z^{d}+z^{-d}$. See [Sil07, Chapter 6] and [Mil06b].
@@
These definitions depend only on the conjugacy class of $f$. Moreover, it is well-known that $f$ is exceptional if and only if some iterate (or every iterate) $f^{\circ k}$ is exceptional ($k\in\mathbb{Z}_{>0}$).
Fix an integer $d\geq 2$. Let $\mathrm{Rat}_{d}$ be the space of degree $d$ endomorphisms on $\mathbb{P}^{1}$, which is a smooth affine variety of dimension $2d+1$ [Sil12]. Let $\mathrm{FL}_{d}\subseteq\mathrm{Rat}_{d}$ be the locus of flexible Lattès maps, which is Zariski closed in $\mathrm{Rat}_{d}$. The group $\mathrm{PGL}_{2}=\mathrm{Aut}(\mathbb{P}^{1})$ acts on $\mathrm{Rat}_{d}$ by conjugacy. The geometric quotient
$\mathcal{M}_{d}:=\mathrm{Rat}_{d}/\mathrm{PGL}_{2}=\mathrm{Spec}(\mathcal{O}(\mathrm{Rat}_{d})^{\mathrm{PGL}_{2}})$
is the (coarse) moduli space of endomorphisms of degree $d$ [Sil12], which is an affine variety of dimension 2d22d-2 [Sil07, Theorem 4.36(c)]. Let $\Psi:\mathrm{Rat}_{d}\to\mathcal{M}_{d}$ be the quotient morphism. For $f\in\mathrm{Rat}_{d}(\mathbb{C})$, we also denote its conjugacy class by $\Psi(f)=[f]\in\mathcal{M}_{d}(\mathbb{C})$. Set $\mathrm{FL}_{d}]:=\Psi(\mathrm{FL}_{d})$. Note that $\mathrm{FL}_{d}$ is non-empty if and only if $d$ is a perfect square in $\mathbb{Z}$, and in this case $\mathrm{FL}_{d}]$ is equidimensional of dimension $1$. The above construction works over any algebraically closed field of characteristic zero and commutes with base changes.
Fix $n\in\mathbb{Z}_{>0}$. Let $\mathrm{Per}_{n}(f)=\mathrm{Per}_{n}(f)(\mathbb{C})$ be the multiset  [^1] A multiset is a collection that allows multiple instances for each of its elements. The number of instances of an element is called its multiplicity. For example, $\{a,a,b,c,c,c\}$ is a multiset of cardinality $6$, with multiplicities $2,1,3$ for $a,b,c$, respectively. of fixed points of $f^{\circ n}$. This multiset has cardinality $d^{n}+1=:N_{d,n}$. Using elementary symmetric polynomials, the multipliers at the $N_{d,n}$ fixed points of $f^{\circ n}$ define a point
@@
where $\sigma_{j,n}(f)$ is the $j$-th elementary symmetric polynomial in the $N_{d,n}$-tuple $\{\rho_{f^{\circ n}}(z)\}_{z\in\mathrm{Per}_{n}(f)}$ for $1\leq j\leq N_{d,n}$. We define the multiplier spectrum of $f$ to be the sequence $S(f)=(S_{n}(f))_{n=1}^{\infty}$. It is clear that $S_{n}$ takes the same value within a conjugacy class of rational maps and thus defines a morphism $S_{n}:\mathcal{M}_{d}\to\mathbb{A}^{N_{d,n}}$ over $\mathbb{C}$, called the multiplier spectrum morphism of level $n$. (In fact, the morphism $S_{n}$ is defined over $\mathbb{Q}$.) The morphism
$\tau_{d,n}:\mathcal{M}_{d}\to\mathbb{A}^{N_{d,1}}\times\cdots\times\mathbb{A}^{N_{d,n}}=\mathbb{A}^{N_{d,1}+\cdots+N_{d,n}}$
given by $\tau_{d,n}([f])=\left(S_{1}([f]),\dots,S_{n}([f])\right)$ is called the multiplier spectrum morphism up to level $n$. For $f\in\mathrm{Rat}_{d}(\mathbb{C})$, we also write $\tau_{d,n}(f):=\tau_{d,n}([f])$.
For $n\in\mathbb{Z}_{>0}$, set $R_{d,n}:=\{([f],[g])\in\mathcal{M}_{d}(\mathbb{C})^{2}\mid\tau_{d,n}(f)=\tau_{d,n}(g)\}$. Then $(R_{d,n})_{n=1}^{\infty}$ forms a decreasing sequence of Zariski closed subsets of $\mathcal{M}_{d}(\mathbb{C})^{2}$, which stabilizes due to the Noetherian property. Let $m_{d}\geq 1$ be minimal such that $\cap_{n=1}^{\infty}R_{d,n}=R_{d,m_{d}}$. Consequently, for all $g\in\mathrm{Rat}_{d}(\mathbb{C}), ff$ and $g$ have the same multiplier spectrum $S(f)=S(g)$ if and only if $\tau_{d,m_{d}}(f)=\tau_{d,m_{d}}(g)$, i.e., $S_{n}(f)=S_{n}(g)$ for $1\leq n\leq m_{d}$. Define $\tau_{d}:=\tau_{d,m_{d}}:\mathcal{M}_{d}\to\mathbb{A}^{N_{d,1}+\cdots+N_{d,m_{d}}}$, called the multiplier spectrum morphism on $\mathcal{M}_{d}$. Abusing notation, we also use $\tau_{d}$ to denote $\tau_{d}\circ\Psi$, called the multiplier spectrum morphism on $\mathrm{Rat}_{d}$. For two rational maps $g\in\mathrm{Rat}_{d}(\mathbb{C}), ff$ and $g$ are called isospectral if they have the same multiplier spectrum $S(f)=S(g)$ (or equivalently, if $\tau_{d}(f)=\tau_{d}(g)$). It is well-known that rational maps within an irreducible component of $\mathrm{FL}_{d}(\mathbb{C})$ are isospectral.
@@
$\tau_{d}:\mathcal{M}_{d}(\mathbb{C})\setminus\Psi(\mathrm{FL}_{d}(\mathbb{C}))\to\mathbb{A}^{N_{d,1}}(\mathbb{C})\times\cdots\times\mathbb{A}^{N_{d,m_{d}}}(\mathbb{C})$
@@
Similarly, by replacing multipliers with lengths and $\mathbb{C}$ with $\mathbb{R}$ in the above definitions, we obtain for each $n\in\mathbb{Z}_{>0}$ the length spectrum map $L_{n}:\mathcal{M}_{d}(\mathbb{C})\to\mathbb{R}^{N_{d,n}}$ of level $n$. The length spectrum of $f$ is defined as the sequence $L(f)=(L_{n}(f))_{n=1}^{\infty}$. We define the length spectrum map up to level $n$ as
$\eta_{d,n}:\mathcal{M}_{d}(\mathbb{C})\to\mathbb{R}^{N_{d,1}}\times\cdots\times\mathbb{R}^{N_{d,n}},[f]\mapsto(L_{1}(f),\dots,L_{n}(f)).$
(For the time being, we treat these maps merely as set-theoretic maps.)
@@ 1.2. Statement of the main results
Fix an integer $d\geq 2$. By restriction, we obtain the multiplier spectrum morphism for polynomials. Let $\mathrm{Poly}^{d}\subset\mathrm{Rat}_{d}$ be the closed subvariety of polynomials of degree $d$, and let $\mathrm{MPoly}^{d}:=\mathrm{Poly}^{d}/\mathrm{Aff}$ be the moduli space of polynomials of degree $d$. See [FG22]. Here $\mathrm{Aff}$ is the group of automorphisms of $\mathbb{A}^{1}$ consisting of all degree one polynomials, acting on $\mathrm{Poly}^{d}$ by conjugacy. So $\mathrm{Aff}(\mathbb{C})=\{az+b\in\mathbb{C}[z]\colon a\in\mathbb{C}^{\times},b\in\mathbb{C}\}$. The following conditions are equivalent for all polynomials $f,g\in\mathrm{Poly}^{d}(\mathbb{C})$:
  • there exists $\tau\in\mathrm{PGL}_{2}(\mathbb{C})$ such that $f=\tau\circ g\circ\tau^{-1}$;
  • there exists $\sigma\in\mathrm{Aff}(\mathbb{C})$ such that $f=\sigma\circ g\circ\sigma^{-1}$;
This allows us to view $\mathrm{MPoly}^{d}\subset\mathcal{M}_{d}$ as a subset of $\mathcal{M}_{d}$. We abuse notation and let $\Psi:\mathrm{Poly}^{d}\to\mathrm{MPoly}^{d}$ still denote the quotient morphism. For every $f\in\mathrm{Poly}^{d}(\mathbb{C})$, we also use $\Psi(f)=[f]\in\mathrm{MPoly}^{d}(\mathbb{C})$ to denote the (affine) conjugacy class of $f$ in $\mathrm{Poly}^{d}(\mathbb{C})$. The inclusions $\mathrm{MPoly}^{d}\stackrel{{\scriptstyle\iota_{0}}}{{\hookrightarrow}}\mathcal{M}_{d}$ and $\mathrm{Poly}^{d}\stackrel{{\scriptstyle\iota}}{{\hookrightarrow}}\mathrm{Rat}_{d}$ are compatible with $\Psi$, i.e., $\Psi\circ\iota=\iota_{0}\circ\Psi$. For every integer $n\geq 1$, denote the restriction of $\tau_{d,n}$ (resp. $\tau_{d}$) to the moduli space $\mathrm{MPoly}^{d}$ of polynomials by $\tilde{\tau}_{d,n}$ (resp. $\tilde{\tau}_{d}$). Similarly, we obtain $\tilde{\eta}_{d,n}$ for the length spectrum.
@@
Ji and Xie also proved the following polynomial version of Theorem 1.3, which was also proved independently by Huguin [Hug24] using completely different methods:
@@
$\tilde{\tau}_{d}:\mathrm{MPoly}^{d}(\mathbb{C})\to\mathbb{A}^{N_{d,1}}(\mathbb{C})\times\cdots\times\mathbb{A}^{N_{d,m_{d}}}(\mathbb{C})$
@@
The proof of Theorem 1.4 in [JX25] parallels that of Theorem 1.3 in the same paper. Specifically, the proof of Theorem 1.3 relies on two key ingredients:
  • a DAO-type theorem [JX25, Theorem 3.4] for curves (here “DAO” means “Dynamical André-Oort”);
  • a result [JX25, Theorem 3.3] about “general” rational maps, which is based on Pakovich’s work [Pak25a].
In contrast, the proof of Theorem 1.4 in [JX25] just replaces the second ingredient [JX25, Theorem 3.3] with [FG22, Theorems 3.51 and 3.52]. Huguin’s proof [Hug24] of Theorem 1.4 actually shows that $\tilde{\tau}_{d,2}$ is generically injective for every $d\geq 2$, and it is based on intricate local computations.
The class of simple rational maps is involved in the second ingredient [JX25, Theorem 3.3]. A rational map $G\in\mathbb{C}(z)$ of degree $d\geq 2$ is called simple if $G$ has exactly 2d22d-2 distinct critical values in $\mathbb{P}^{1}(\mathbb{C})$. However, a polynomial of degree $d\geq 3$ is never simple, due to the totally invariant fixed point $\infty$. To overcome this problem, for a polynomial $F\in\mathbb{C}[z]$ of degree $d\geq 2$, we call $F$ pre-simple if it has exactly d1d-1 distinct critical values in $\mathbb{C}$. This class of pre-simple polynomials serves as an analogue of simple rational maps for the polynomial case. We observe that Pakovich’s methods [Pak25a] are applicable to pre-simple polynomials. In § 3.1, we present the following polynomial version of [JX25, Theorem 3.3] (for the case $d\geq 4$), which is essentially due to Pakovich [Pak25a]:
@@
For every integer $d\geq 4$, there exists a non-empty Zariski open subset $U$ of $\mathrm{Poly}^{d}(\mathbb{C})$ such that, for every $f\in U$, the following hold:
@@
    for every pre-simple $g\in\mathrm{Poly}^{d}(\mathbb{C})$, $f$ and $g$ are intertwined if and only if $[f]=[g]$ in $\mathrm{MPoly}^{d}(\mathbb{C})$.
Here, two polynomials $F,G\in\mathbb{C}[z]$ of degree $\geq 2$ are called intertwined if there exists a (possibly reducible) algebraic curve $Z\subset\mathbb{A}^{2}$ whose projections to both axes are onto such that $Z$ is invariant under the endomorphism
$(F,G):\mathbb{A}^{2}\to\mathbb{A}^{2},(x,y)\mapsto(F(x),G(y)).$
Remark 1.6.
Let $F,G\in\mathbb{C}[z]$ be polynomials of degree $\geq 2$. It is known that $F$ and $G$ are intertwined if and only if there exist non-constant polynomials $R,h_{1},h_{2}\in\mathbb{C}[z]\setminus\mathbb{C}$ and $n\in\mathbb{Z}_{>0}$ such that
$F^{\circ n}\circ h_{1}=h_{1}\circ R\quad\text{and}\quad G^{\circ n}\circ h_{2}=h_{2}\circ R.$
In particular, $\deg(F)=\deg(G)$. See [FG22, Theorem 3.39].
Recently, Pakovich [Pak25b] studied general rational maps of degree $2\leq d\leq 3$ in detail. Consequently, one could get a proof of Theorem 1.3 for this case, using the results in [Pak25b] as part of the argument. The philosophy is that a general rational map of degree $d$ has “good” properties regarding the decompositions of its iterates. However, the simplicity condition (used in [Pak25a] where $d\geq 4$) is not sufficient for the cases $2\leq d\leq 3$.
In § 3.2, following Pakovich [Pak25b], we study general polynomial maps of degree $2\leq d\leq 3$ and obtain the following polynomial version of [JX25, Theorem 3.3] for these lower degree cases: